# Underdetermined Fourier Extensions for Surface Partial Differential Equations

### v2

@@ 1. Introduction
Among existing approaches, surface finite element methods are likely the best understood, owing largely to the significant work of Dziuk and Elliott in the 2010s [dziuk13]. A number of implementations of radial basis function (RBF) methods, often using least squares collocation, have also been studied. RBFs are well understood for interpolation and can be high order, and while the popular Kansa’s [kansa90] is known to potentially fail, techniques based on oversampling have been successful on flat domains as well as surfaces when a point cloud regularity condition is satisfied [chen20]. The Closest Point Method (CPM) [ruuth08] has been extensively studied numerically, and consistency follows directly from consistency of the underlying interpolation and finite difference techniques, but stability is not yet fully understood, though the method works in practice in most cases. RBF methods using similar extension approaches as the CPM, either by performing a Closest Point extension directly [petra18], or by satisfying normal derivative conditions [piret12], have also been proposed.
@@
Norm-minimizing Hermite-Birkhoff interpolants have been previously studied in the RBF literature through Hermite RBFs [frank98, sun94]. More recently, Chandrasekaran, Gorman, and Mhaskar have studied solving the norm-minimization problem more directly [chand18], which has some numerical advantages. Namely, solving the optimization problem directly can improve conditioning and is often simpler to implement. In our work, we explore the direct solution of the norm-minimization problem for surface PDEs via function extension from surfaces to a box. Extension does not require a basis for functions on the surface; a standard Fourier basis can be used.
Along with numerical tests in Section 4, we present new analytical results in Propositions 9, 10, and 11 showing the high-order nature of our extension approach. We also prove a very general result regarding the relationship between boundedness, PDE solvability, and convergence in Proposition 12. In particular, we show that boundedness of numerical solutions to a PDE as point spacing goes to zero implies solvability of the non-discretized PDE and convergence of the numerical solution to a PDE solution, regardless of whether the problem has multiple solutions. This result is used to find eigenvalues in Subsection 4.2.
@@ 2.1. Fourier Extensions
A possible resolution to this problem is to avoid finding eigenfunctions on a complicated domain altogether. Instead, the possibly complicated domain $S$ is placed inside a box-shaped domain $\Omega\subseteq\mathbb{R}^{m}$. We can then use the box’s Fourier basis functions to expand functions on $S$, albeit not uniquely. In approximation theory terms, our functions $\left\{\phi_{n}\right\}_{n=1}^{N_{b}}$ will not form a basis for functions on $S$, but rather a frame [adcoc19].
@@
This sort of problem, where we extend $f$ from $S$ to a larger domain $\Omega\supset S$ using a Fourier series, is referred to by Boyd [boyd03] as Fourier extension of the third kind when $S$ and $\Omega$ have the same dimensionality. One motivation for computing such an extension is that we are able to compute derivatives of Fourier series on $\Omega$ with high accuracy.
Computing a Fourier extension without knowledge of the function’s exact Fourier coefficients is typically done by only approximately satisfying $f=\tilde{f}$ on $S$. A sample of points in $S$ is chosen, then the coefficients $a_{n}$ are computed via least squares. Certain implementations have been shown to converge super-algebraically in 1D for smooth functions when using Chebyshev nodes, or oversampled, uniformly spaced nodes [adcoc14].
Fourier extensions have been applied successfully in various contexts, ranging from straightforward 1D function approximation, to surface reconstruction [bruno07], to PDEs in flat domains [boyd05, stein17, bruno22, matth18]. Analytical results, however, remain fairly sparse in higher dimensions, particularly for PDEs.
@@ 2.2. Closest Point and Closest Point-like Extensions
A straightforward method for extending functions off of manifolds is a Closest Point extension, which is the idea underlying the Closest Point Method (CPM) for PDEs on surfaces [ruuth08].
@@
The (Euclidean) Closest Point extension of a function $f:S\to\mathbb{R}$, is the function $f\circ\text{cp}$, where
$\text{cp}\left(\boldsymbol{x}\right)=\underset{\boldsymbol{y}\in S}{\text{argmin}}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|_{2}.$
If $S$ is a twice differentiable manifold, then cp is well-defined and differentiable in a neighbourhood of an open subset of $S$. The utility of such an extension is clear from the following facts (see [ruuth08] for the original statements and [marz12] for detailed proofs).
@@
$\nabla\left(f\circ\text{cp}\right)\biggr{|}_{S}=\nabla_{S}f,$
@@
$\Delta\left(f\circ\text{cp}\right)\biggr{|}_{S}=\Delta_{S}f,$
@@
As it turns out, Closest Point extensions are not the only extensions where Propositions 1 and 2 hold. März and Macdonald studied extensions of the form $f\circ P$, where $P:\Omega\to S$ is idempotent, and found general conditions needed for $P$ so that Propositions 1 and 2 still hold with $P$ instead of cp [marz12].
We state a couple of facts needed for functions from $\Omega\to S$ in general to be used for computing surface differential operators. First, we need a definition.
@@
Let $f\in C^{1}\left(S\right)$ where $S$ is a twice differentiable manifold. Then $\tilde{f}\in C^{1}\left(\Omega\right)$ is a first-order Closest Point-like extension of $f$ when
@@
To work with the Laplace-Beltrami operator, we need an important result adapted from Xu and Zhao [xu03].
@@
[xu03, Lemma 1] Let $f\in C^{2}\left(S\right),\tilde{f}\in C^{2}\left(\Omega\right)$ where $S$ is a thrice differentiable manifold. If
@@
Crucially, the only surface information we need is $\hat{\boldsymbol{n}}_{S}$, not any of its derivatives or curvature information. Also note that Corollary 3 only imposes one additional interpolation condition. Some previous methods, such as the higher-order version of Piret’s orthogonal gradients method with RBFs [piret12], impose two additional conditions; $\hat{\boldsymbol{n}}_{s}\cdot\nabla_{S}f$ and $\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{f}\right)\hat{\boldsymbol{n}}_{S}$ are both set to zero. Using only one condition is more computationally efficient. We use this corollary to solve surface PDEs in Section 4.
We now introduce one more fact and a definition before another preliminary result.
@@
For every $C^{p}$ manifold $S\subset\mathbb{R}^{n}$, for small enough $\beta>0$, where $S\subset U_{\beta}$ and $U_{\beta}$ contains all points with distance less than $\beta$ to $S$, there exists a signed distance function $\varphi_{\beta}\in C^{p}\left(U_{\beta}\right)$. That is, $\varphi_{\beta}$ is $C^{p}$ and $\left|\varphi_{\beta}\left(\boldsymbol{x}\right)\right|=\inf_{\boldsymbol{y}\in S}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|_{2}<\beta$.
Proposition 4 is a simple consequence of Lemma 14.16 from [gilba01].
Definition 3.
A $C^{\infty}$ top hat function is a function $\Phi\in C^{\infty}\left(\mathbb{R}\right)$ such that for some $\alpha,\beta>0$, where $\alpha<\beta$,
$\Phi\left(x\right)$ $=1\text{, for $x\in\left[-\alpha,\alpha\right]$}$
$\Phi\left(x\right)$ $=0\text{, for $x\notin\left[-\beta,\beta\right]$}.$
There are many examples of smooth top hat functions [boyd06]. Using the Euclidean Closest Point function and top hat functions, we can now show that smooth, periodic, Closest Point-like extensions exist.
Proposition 5.
For $p\geq 1$, $f\in C^{p}\left(S\right)$, and a $C^{p+1}$ manifold $S$, there exists a periodic function $\tilde{f}\in C^{p}\left(\Omega\right)$ such that $\tilde{f}$ is a first-order Closest Point-like extension of $f$.
This is constructed as $\left(f\circ\text{cp}\right)\cdot\left(\Phi\circ\varphi_{\beta}\right)$ within a distance $\beta$ of $S$ and zero elsewhere, where $\varphi_{\beta}$ is as in Proposition 4 and $\Phi$ is a smooth top hat function with compact support of width $2\beta$ as in Definition 3. Note that $\text{cp}\left(\boldsymbol{x}\right)=\boldsymbol{x}-\varphi_{\beta}\left(\boldsymbol{x}\right)\nabla\varphi_{\beta}\left(\boldsymbol{x}\right)$, which is why $S$ must be $C^{p+1}$. This result tells us that for the PDEs on surfaces we consider, there are periodic extended solutions with the same degree of smoothness as the solution on the surface. In particular, the constrained norm-minimization problems that we set up will have feasible solutions.
@@ 2.3. Radial Basis Function Interpolation
Radial basis functions (RBFs) are a class of functions used for interpolation and numerical PDEs that depend on the location of the interpolation points $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset S\subset\mathbb{R}^{n}$. This is vital since, due to the Mairhuber-Curtis Theorem [curti59, mairh56], any set of $\tilde{N}$ functions that does not depend on the location of interpolation points cannot uniquely interpolate every set of $\tilde{N}$ points. Specifically, polynomial interpolation when the number of polynomials matches the number of points, can fail to produce a solution.
@@
$h_{\text{max}}=\max\left(\left\{\frac{1}{2}\left|x_{k}-x_{k-1}\right|\right\}_{k=2}^{\tilde{N}}\cup\left\{x_{1}-a,b-x_{\tilde{N}}\right\}\right).$
@@
A positive definite radial basis function is a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ that posesses two properties:
@@
First, we need a definition (see Definition 3.6 of [wendl04]) to restrict the types of domains we can consider.
@@
A set $\Omega\subset\mathbb{R}^{m}$ satisfies an interior cone condition if there exists an angle $\theta\in\left(0,\frac{\pi}{2}\right)$ and a radius $r>0$ such that for all $\boldsymbol{x}\in\Omega$, there exists some unit vector $\boldsymbol{\xi}\left(\boldsymbol{x}\right)$ so that
$\left\{\boldsymbol{x}+\lambda\boldsymbol{y}:\boldsymbol{y}\in\mathbb{R}^{d},\left\|\boldsymbol{y}\right\|_{2}=1,\boldsymbol{y}\cdot\boldsymbol{\xi}\left(\boldsymbol{x}\right)\geq\cos\theta,\lambda\in\left[0,r\right]\right\}\subseteq\Omega.$
In other words, $\Omega$ satisfies an interior cone condition if there is some cone of a fixed size and angle that can be placed with its vertex at each $\boldsymbol{x}\in\Omega$ and remain entirely contained in $\Omega$. Informally, $\Omega$ is only “finitely pointy”.
The next proposition is adapted from Wendland [wendl04].
@@
[wendl04, Thms. 10.35 and 11.17] Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(\tilde{\Omega}\right)$ with $m\geq 3$ if $q=0$ and $m\in\mathbb{N}$ for $q>0$ where $\tilde{\Omega}\supset\Omega$ is bounded and $\Omega\subset\mathbb{R}^{m}$ satisfies an interior cone condition. Then there exists a radial basis function $\phi_{m,q}$ such that if $\tilde{f}$ is the RBF interpolant of $f$ on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ using $\phi_{m,q}$ where $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ has fill distance $h_{\text{max}}$ on $\Omega$, then there exist constants $C_{m,q},h_{0,m,q}>0$ such that as long as $h_{\text{max}}\leq h_{0,m,q}$,
$\left\|f-\tilde{f}\right\|_{L^{\infty}\left(\Omega\right)}\leq C_{m,q}h_{\text{max}}^{q+\frac{1}{2}}\left\|f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(\Omega\right)}.$
@@
Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(\tilde{\Omega}\right)$ with $m\geq 3$ if $q=0$ and $m\in\mathbb{N}$ for $q>0$ where $\tilde{\Omega}\supset\Omega$ is bounded and $\Omega$ satisfies an interior cone condition. Assume $f$ vanishes on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ with fill distance $h_{\text{max}}$ on $\Omega$, then for all $p\in\left\{0,1,\ldots q\right\}$ for $m\geq 3$ and $p\in\left\{1,\ldots q\right\}$ for $m<3$ there exist constants $C_{m,q},h_{0,m,q}>0$ such that as long as $h_{\text{max}}\leq h_{0,m,q}$,
$\left\|f\right\|_{L^{\infty}\left(\Omega\right)}\leq C_{m,q}h_{\text{max}}^{p+\frac{1}{2}}\left\|f\right\|_{H^{\frac{m}{2}+p+\frac{1}{2}}\left(\Omega\right)}.$
Similar results regarding error bounds for functions with scattered zeros exist in the literature, and could also be used for efficient convergence proofs for a wide variety of methods, including our own. In particular, more general statements for convergence in Sobolev norms have been shown by Narcowich, Ward, and Wendland [narco05].

### v3

@@ 1. Introduction
Among existing approaches, surface finite element methods are likely the best understood, owing largely to the significant work of Dziuk and Elliott in the 2010s (see Dziuk and Elliott, 2013). A number of implementations of radial basis function (RBF) methods, often using least squares collocation, have also been studied. RBFs are well understood for interpolation and can be high order, and while the popular method of Kansa (1990) is known to potentially fail, techniques based on oversampling have been successful on flat domains as well as surfaces when a point cloud regularity condition is satisfied (Chen and Ling, 2020). The Closest Point Method (CPM) (Ruuth and Merriman, 2008) has been extensively studied numerically, and consistency follows directly from consistency of the underlying interpolation and finite difference techniques, but stability is not yet fully understood, though the method works in practice in most cases. RBF methods using similar extension approaches as the CPM, either by performing a Closest Point extension directly (Petras et al., 2018), or by satisfying normal derivative conditions (Piret, 2012), have also been proposed.
@@
Norm-minimizing Hermite-Birkhoff interpolants have been previously studied in the RBF literature through Hermite RBFs (see, for example, Franke and Schaback, 1998; Sun, 1994). More recently, Chandrasekaran, Gorman, and Mhaskar have studied solving the norm-minimization problem more directly (Chandrasekaran et al., 2018), which has some numerical advantages. Namely, solving the optimization problem directly can improve conditioning and is often simpler to implement. In our work, we explore the direct solution of the norm-minimization problem for surface PDEs via function extension from surfaces to a box. Extension does not require a basis for functions on the surface; a standard Fourier basis can be used.
Along with numerical tests in Section 4, we present new analytical results in Propositions 8, 9, and 11 showing the high-order nature of our extension approach. We also prove a very general result regarding the relationship between boundedness, PDE solvability, and convergence in Proposition 10. In particular, we show that boundedness of numerical solutions to a PDE as point spacing goes to zero implies solvability of the non-discretized PDE and convergence of the numerical solution to a PDE solution, regardless of whether the problem has multiple solutions. This result is used to find eigenvalues in Subsection 4.2.
@@ 2.1. Fourier Extensions
A possible resolution to this problem is to avoid finding eigenfunctions on a complicated domain altogether. Instead, the possibly complicated domain $S$ is placed inside a box-shaped domain $\Omega\subseteq\mathbb{R}^{m}$. We can then use the box’s Fourier basis functions to expand functions on $S$, albeit not uniquely. In approximation theory terms, our functions $\left\{\phi_{n}\right\}_{n=1}^{N_{b}}$ will not form a basis for functions on $S$, but rather a frame (see Adcock and Huybrechs, 2019).
@@
This sort of problem, where we extend $f$ from $S$ to a larger domain $\Omega\supset S$ using a Fourier series, is referred to by Boyd (2002) as Fourier extension of the third kind when $S$ and $\Omega$ have the same dimensionality. One motivation for computing such an extension is that we are able to compute derivatives of Fourier series on $\Omega$ with high accuracy.
Computing a Fourier extension without knowledge of the function’s exact Fourier coefficients is typically done by only approximately satisfying $f=\tilde{f}$ on $S$. A sample of points in $S$ is chosen, then the coefficients $a_{n}$ are computed via least squares. Certain implementations have been shown to converge super-algebraically in 1D for smooth functions when using Chebyshev nodes, or oversampled, uniformly spaced nodes (Adcock et al., 2014).
Fourier extensions have been applied successfully in various contexts, ranging from straightforward 1D function approximation, to surface reconstruction (Bruno et al., 2007), to PDEs in flat domains (Boyd, 2005; Stein et al., 2017; Bruno and Paul, 2022; Matthysen and Huybrechs, 2018). Analytical results, however, remain fairly sparse in higher dimensions, particularly for PDEs.
@@ 2.2. Closest Point and Closest Point-like Extensions
A straightforward method for extending functions off of manifolds is a Closest Point extension, which is the idea underlying the Closest Point Method (CPM) for PDEs on surfaces (Ruuth and Merriman, 2008).
@@
The (Euclidean) Closest Point extension of a function $f:S\to\mathbb{R}$, is the function $f\circ\text{cp}_{S}$, where
$\text{cp}_{S}\left(\boldsymbol{x}\right)=\underset{\boldsymbol{y}\in S}{\text{argmin}}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|_{2}.$
If $S$ is a twice differentiable manifold, then $\text{cp}_{S}$ is well-defined and differentiable on an open set in $\mathbb{R}^{m}$ containing $S$ (see, for example, Lee, 2003, Problem 6-5); $\text{cp}_{S}$ in general is one order less smooth than the manifold. The utility of such an extension is clear from the following facts (see the work of Ruuth and Merriman (2008) for the original statements and März and Macdonald (2012) for detailed proofs).
@@
$\nabla\left(f\circ\text{cp}_{S}\right)\biggr{|}_{S}=\nabla_{S}f,$
@@
$\Delta\left(f\circ\text{cp}_{S}\right)\biggr{|}_{S}=\Delta_{S}f,$
@@
As it turns out, Closest Point extensions are not the only extensions where Facts 1 and 2 hold. März and Macdonald (2012) studied extensions of the form $f\circ P$, where $P:\Omega\to S$ is idempotent, and found general conditions needed for $P$ so that Facts 1 and 2 still hold with $P$ instead of $\text{cp}_{S}$.
We state a couple of facts needed for functions extended from $S$ to $\Omega$ in general to be used for computing differential operators on a hypersurface $S$. First, we need a definition.
@@
Let $f\in C^{1}\left(S\right)$ where $S$ is a twice differentiable hypersurface. Then $\tilde{f}\in C^{1}\left(\Omega\right)$ is a first-order Closest Point-like extension of $f$ from $S$ when
@@
To work with the Laplace-Beltrami operator, we need an important result adapted from Xu and Zhao (2003).
@@
(Xu and Zhao, 2003, Lemma 1) Let $f\in C^{2}\left(S\right),\tilde{f}\in C^{2}\left(\Omega\right)$ where $S$ is a thrice differentiable hypersurface. If
@@
Crucially, the only hypersurface information we need is $\hat{\boldsymbol{n}}_{S}$, not any of its derivatives or curvature information. Also note that Corollary 3 only imposes one additional interpolation condition. Some previous methods, such as the higher-order version of Piret’s orthogonal gradients method with RBFs (Piret, 2012), impose two additional conditions; $\hat{\boldsymbol{n}}_{s}\cdot\nabla_{S}f$ and $\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{f}\right)\hat{\boldsymbol{n}}_{S}$ are both set to zero. Using only one condition is more computationally efficient. We use this corollary to solve surface PDEs in Section 4.
We now note that smooth, periodic, Closest Point-like extensions exist for suitable hypersurfaces and functions.
@@
Let $S\subset\Omega$ be a $C^{p+1}$ hypersurface such that $\overline{S}\subset T$, where $T\subset\Omega$ is also a $C^{p+1}$ hypersurface. For $p\geq 1$, $f\in C^{p}\left(T\right)$, there exists a periodic function with $p$ periodic derivatives $\tilde{f}\in C^{p}\left(\Omega\right)$ such that $\tilde{f}$ is a first-order Closest Point-like extension of $f$ from $S$.
This is constructed as $\tilde{f}=\left(f\circ\text{cp}_{T}\right)$ in an open set $U$ containing $T$, then extended from a closed subset of $U$ with an interior that contains $\overline{S}$ to $\Omega$ using a partition of unity as in Lemma 2.26 of the text by Lee (2003), such that the support of $\tilde{f}$ is compactly contained in $\Omega$. This result tells us that for the PDEs on surfaces we consider, there are periodic extended solutions with the same degree of smoothness as the solution on the surface. In particular, the constrained norm-minimization problems that we set up will have feasible solutions.
@@ 2.3. Radial Basis Function Interpolation
Radial basis functions (RBFs) are a class of functions used for interpolation and numerical PDEs that depend on the location of the interpolation points $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset S\subset\mathbb{R}^{n}$. This is vital since, due to the Mairhuber-Curtis Theorem (Curtis Jr., 1959; Mairhuber, 1956), any set of $\tilde{N}$ functions that does not depend on the location of interpolation points cannot uniquely interpolate every set of $\tilde{N}$ points. Specifically, polynomial interpolation when the number of polynomials matches the number of points can fail to produce a solution.
@@
[…]
@@
A positive definite radial basis function is a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ that possesses two properties:
@@
First, we need a definition (see Wendland, 2004, Def. 3.6) to restrict the types of domains we can consider.
@@
A set $U\subset\mathbb{R}^{m}$ satisfies an interior cone condition if there exists an angle $\theta\in\left(0,\frac{\pi}{2}\right)$ and a radius $r>0$ such that for all $\boldsymbol{x}\in U$, there exists some unit vector $\boldsymbol{\xi}\left(\boldsymbol{x}\right)$ so that
$\left\{\boldsymbol{x}+\lambda\boldsymbol{y}:\boldsymbol{y}\in\mathbb{R}^{m},\left\|\boldsymbol{y}\right\|_{2}=1,\boldsymbol{y}\cdot\boldsymbol{\xi}\left(\boldsymbol{x}\right)\geq\cos\theta,\lambda\in\left[0,r\right]\right\}\subseteq U.$
In other words, $U$ satisfies an interior cone condition if there is some cone of a fixed size and angle that can be placed with its vertex at each $\boldsymbol{x}\in U$ and remain entirely contained in $U$. Informally, $U$ is only “finitely pointy”. Hypersurfaces do not satisfy an interior cone condition; the hypersurfaces we consider will require that the domain $U$ of a local chart satisfies an interior cone condition instead.
The next proposition is adapted from Wendland (2004).
@@
(Wendland, 2004, Thms. 10.35 and 11.17) Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(U\right)$ with $m\geq 3$ if $q=0$ and $m\in\mathbb{N}$ for $q>0$ where $U\subset\mathbb{R}^{m}$ is bounded and satisfies an interior cone condition. Then there exists a radial basis function $\phi_{m,q}$ such that if $\tilde{f}$ is the RBF interpolant of $f$ on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ using $\phi_{m,q}$ where $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ has fill distance $h_{\text{max}}$ on $U$, then there exist constants $C_{m,q,\left|\alpha\right|},h_{0,m,q,\left|\alpha\right|}>0$ such that as long as $h_{\text{max}}\leq h_{0,m,q,\left|\alpha\right|}$,
$\left\|\partial^{\alpha}f-\partial^{\alpha}\tilde{f}\right\|_{L^{\infty}\left(U\right)}\leq C_{m,q,\left|\alpha\right|}h_{\text{max}}^{q+\frac{1}{2}-\left|\alpha\right|}\left\|f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(U\right)},$
where $\alpha$ is a multi-index with $\left|\alpha\right|\in\left\{0,1,2,\ldots,q\right\}$.
@@
Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(U\right)$ with $m\geq 3$ if $q=0$ and $m\in\mathbb{N}$ for $q>0$ where $U$ is bounded and satisfies an interior cone condition. Assume $f$ vanishes on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ with fill distance $h_{\text{max}}$ on $U$, then for all $p\in\left\{0,1,\ldots q\right\}$ for $m\geq 3$ and $p\in\left\{1,2,\ldots q\right\}$ for $m<3$, there exist constants $C_{m,p,\left|\alpha\right|},h_{0,m,p,\left|\alpha\right|}>0$ such that as long as $h_{\text{max}}\leq h_{0,m,p,\left|\alpha\right|}$,
$\left\|\partial^{\alpha}f\right\|_{L^{\infty}\left(U\right)}\leq C_{m,p,\left|\alpha\right|}h_{\text{max}}^{p+\frac{1}{2}-\left|\alpha\right|}\left\|f\right\|_{H^{\frac{m}{2}+p+\frac{1}{2}}\left(U\right)}.$
where $\alpha$ is a multi-index with $\left|\alpha\right|\in\left\{0,1,2,\ldots,p\right\}$.
Similar results regarding error bounds for functions with scattered zeros exist in the literature and could also be used for efficient convergence proofs for a wide variety of methods, including our own. In particular, more general statements for convergence in Sobolev norms have been shown by Narcowich et al. (2005).